Determine x for 10^(2x+1) = 100^x
a. x = -1
b. x = 2
c. x = 0
d. No solution
After working a, b, c....I picked answer d.

Answer by stanbon(48569)

Determine x for 10^(2x+1) = 100^x
a. x = -1
b. x = 2
c. x = 0
d. No solution
--------------------------
10^(2x+1) = 10^2x
2x+1 = 2x
1 = 0
Contradiction:
Ans: No solution
=====================
Cheers,
Stan H.